OpenStudy (anonymous):

Which mixed number falls between three and two-thirds and four and five-sixths ? Answer four and eleven-twelfths three and three-eighths four and one-fourth three and one-half

OpenStudy (anonymous):

first one is bigger than four and five-sixths right? because \(\frac{11}{12}>\frac{5}{6}\)

OpenStudy (anonymous):

three and three-eighths is smaller than three and two-thirds because \(\frac{3}{8}<\frac{2}{3}\)

OpenStudy (anonymous):

how about the next one?

OpenStudy (anonymous):

Is it the last one?

OpenStudy (anonymous):

no, three and a half is less than three and two thirds, because one half is less than two thirds

OpenStudy (anonymous):

C is always the safest bet, and in this case it is right

OpenStudy (anonymous):

ok. :)